AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 320     Entry time: Mon Feb 28 03:16:37 2022
Author: JM, OF 
Subject: Current status 

04:24

Temperature setpoint 110 K

Rates appeared stable throughout the hour ~10^2 peak height. Last 5 minutes rutherford peak height appears reduced, cycling from 10^1 to 10^2. If this continues with sustained low density, temperature will be cycled to 125 K and back down.

Merger OK

Data rate = 1300 kb/sec

Ring avg pressure = 3.1E-10 mbar

Temperatures OK

System checks OK - wr fail attached

Leakage current ~ 3.228 uA

04:33

Rutherford rate appears to have increased and stabilised again around 10^2 peak height.

Attachment 1: 2022-02-28_04-23-06-rates.png  158 kB  | Hide | Hide all
2022-02-28_04-23-06-rates.png
Attachment 2: 2022-02-28_04-22-28-goodEvents.png  103 kB  | Hide | Hide all
2022-02-28_04-22-28-goodEvents.png
Attachment 3: 2022-02-28_04-22-05-correlation8.png  102 kB  | Hide | Hide all
2022-02-28_04-22-05-correlation8.png
Attachment 4: 2022-02-28_04-21-39-ADCitems.png  103 kB  | Hide | Hide all
2022-02-28_04-21-39-ADCitems.png
Attachment 5: 2022-02-28_04-21-17-wr.png  128 kB  | Hide | Hide all
2022-02-28_04-21-17-wr.png
Attachment 6: 20220228_042449.jpg  1.078 MB  | Hide | Hide all
20220228_042449.jpg
ELOG V3.1.4-unknown